OpenStudy (anonymous):

Suppose that f(−4)=−8, g(−4)=5, f′(−4)=5, and g′(−4)=8. Find h′(−4) for the following. 1.) if h(x)=5f(x)−4g(x) find h'(4) 2.) if h(x)=f(x)g(x) find h'(4) 3.) if h(x)=f(x)/g(x) find h'(4) 4.) if h(x)=g(x)/(1+f(x)) find h'(4)

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

So for the first one, what does the derivative of \[h _{(x)} = 5f_{(x)} - 4 g_{(x)}\] look like?

OpenStudy (anonymous):

would i just change f(x) and g(x) to f'(x) and g'(x)

OpenStudy (anonymous):

yup! so then you know values for both f'(x) and g'(x), so just sub those values in

OpenStudy (anonymous):

so the derivative would just be h'(x)=5f'(x)-4g'(x) and then i would plug in 5 for f'(x) and 8 for g'(x)

OpenStudy (anonymous):

I don't get why it says to find h'(-4) for all of them up top but then find h'(4) in the question :/ I think that's what you would do - I thought it was just asking for h'(-4). You technically have no idea about what f(4) and g(4) are, but I'm guessing that's what they're wanting.

OpenStudy (anonymous):

i mistyped its h'(-4)

OpenStudy (anonymous):

phew! Good. then yeah, you got it.

OpenStudy (anonymous):

i got h'(-4)=-7 for the first one

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

ok and i do that for all of them?

OpenStudy (anonymous):

yeah. It's seeing if you know when/how to use the chain rule and such.

OpenStudy (anonymous):

what would the second one be, then?

OpenStudy (anonymous):

-89. i used the product rule

OpenStudy (anonymous):

i will use the quotient rule for the last two

OpenStudy (anonymous):

okay, yeah. Just makin' sure you didn't just switch the f and g to f' and g' :)

OpenStudy (anonymous):

If you want to make your life easier, you can use the product rule for all of them and say h(x) = f(x) (g(x))^-1

OpenStudy (anonymous):

I could never remember the quotient rule :P But whatever makes sense to you is what you should use :)

OpenStudy (anonymous):

for the last one. the derivative of 1+f(x)=f'(x)?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

calculus high five

OpenStudy (anonymous):

ummm sure

OpenStudy (anonymous):

????

OpenStudy (anonymous):

haha, thanks for humoring me

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok i tried them i got the last three wrong

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

i got -7, -89, -89/25, 89/49

OpenStudy (anonymous):

I got zero for the second one, actually

OpenStudy (anonymous):

how

OpenStudy (anonymous):

by being dumb, I'm wrong. one sec

OpenStudy (anonymous):

for the second one you added the two parts you got. You should have subtracted 64 from 25 -- -39

OpenStudy (anonymous):

oh ok. what about the last two

OpenStudy (anonymous):

I got positive 89/25 for the third one

OpenStudy (anonymous):

ok and for the last

OpenStudy (anonymous):

still tryin it. Have another go and see if you set it right :)

OpenStudy (anonymous):

what did you end up with?

OpenStudy (anonymous):

89/49

OpenStudy (anonymous):

The whole formula. I got \[h'_{(x)} = \frac{g'_{(x)}}{(1+f_{(x)})} - \frac{g_{(x)}f'_{(x)}}{(1+f_{(x)})^{2}}\]

OpenStudy (anonymous):

-8/7 - 25/81

OpenStudy (anonymous):

so it would be -81/49. yup its right thanks

OpenStudy (anonymous):

cool :)

OpenStudy (anonymous):

addition derp on my part -8/7 - 25/49